There is evidence that, other things equal, a 10 percent increase in the number of workers having a particular skill level leads to about a 4 percent decline in wages for workers with that skill level. In addition, this 10-to-4 ratio appears to hold true whether the increase in labor supply is caused by domestic changes in labor supply or by an influx of foreign immigrants Instructions: Round your answers to 2 decimal places. a. Suppose that 42,000 computer programmers work in Silicon Valley. If the number of computer programmers in Silicon Valley increases by 5,040 because of a change in US. immigration laws, by how many percentage points would you expect the wage of computer programmers to fall in Silicon Valley? percent b. Suppose that 8,000 full-time cooks work in restaurants In the Denver area. If Denver becomes popular both with U.S. citizens and with foreigners such that 700 full-time cooks move to Denver from other parts of the United States while 540 full-time cooks move to Denver from other countries, by how much would you expect the wages of full time cooks to fall in Denver? percent
Answer #1

a)

The percentage increase in computer programmers = ( 47,040 - 42,000) \div 42,000

The percentage increase in computer programmers = 12%

For a 12% increase in the number of computer programmers, the decrease in the wages of computer programmers = ( 12 x 4 ) \div 10

For a 12% increase in the number of computer programmers, the decrease in the wages of computer programmers = 4.8%

The answer is 4.8%

______________________________________________________________________

b)

The number of full time cooks after migration = 8000 + 700 + 540

The number of full time cooks after migration = 9,240

The percentage increase in the number of cooks = ( 9240 - 800) \div 8000

The percentage increase in the number of cooks = 15.50%

For a 15.5 % increase in the number of full time cooks, the percentage decrease in the wages of cooks = ( 15.5 x 4 ) \div 10

For a 15.5 % increase in the number of full time cooks, the percentage decrease in the wages of cooks = 6.2%

The answer is 6.2%